# Holes in Latent Space: Topological Signatures Under Adversarial Influence

### v1

@@ A.2.1 Discussion of Results
Figures 12 and 13 highlight that:
@@
Interestingly, our findings align with the results in [44], which indicate that memorization tends to emerge in deeper layers where the effective dimensionality shrinks. Consistent with that view, we observe that executed or ignored prompts show a higher dispersion in mid-layers, implying the model invests additional capacity there for those injected instructions. Meanwhile, a refused request is routed into a more compressed region, effectively “shutting down” further representational expansion. In this sense, deeper layers may provide a setting where the network can more sharply discriminate or overfit certain inputs—supporting the idea that final layers reflect a gradually compressed, yet strategically focused representation space.
Figure 12: Layer-wise Dispersion Ratio for Clean vs. Poisoned Examples. The green and red lines depict mean dispersion ratios for clean and poisoned inputs, respectively, at different layer depths. Error bars around each point represent ±$\pm 1$ standard error of the mean (SEM). In early layers (left side), clean data consistently has higher dispersion on average, whereas in mid-layers (center), poisoned surpasses the clean baseline, indicating a re-distribution of representational capacity for the injected prompts. Layers where the difference is statistically significant ($p_{\text{FDR}}<0.05$) are marked with a red asterisk above the higher mean value.
@@ A.3 Cosine Distance of Representations
We analyze the difference representations $\Delta\mathrm{Act}_{\ell}(x_{i})\in\mathbb{R}^{D}$ for corresponding pairs of clean and poisoned inputs. Specifically, for each model and layer, we load up to five pairs of clean and poisoned activation files, compute the difference between the activations for each pair, and concatenate these differences. From these differences, we draw equal-size subsamples of 5,000 vectors. For each layer and comparison condition, we compute the mean pairwise cosine distance within each subsample. Because cosine distance is scale-invariant, we do not normalize these difference representations. We perform four comparison conditions: clean vs. poisoned, clean vs. clean (where clean samples are split in half), poisoned vs. poisoned (where poisoned samples are split in half), and mixed vs. mixed (where two separate mixed subsamples are created, each containing half clean and half poisoned differences). For each comparison, we generate two distributions of mean pairwise intra-class distances (or inter-class in the clean vs poisoned case) using 3 bootstrap iterations. We then apply Welch’s $t$-test to these distributions to assess whether they diverge significantly.
@@ B.1 Pipeline
We use Ripser [5] to compute barcodes, which is based on Vietoris–Rips filtrations (see Section 2.1). The computational constraints of PH make it impossible to compute the barcode of any of our two datasets (clean vs. poisoned or locked vs. elicited). Therefore, we leverage subsampling approaches (e.g., [14]) and compute barcodes from $K=600$ subsamples $\{x_{i_{1},\ell},\,\dots,\,x_{i_{k},\ell}\}\subset\mathbb{R}^{D}$ with size $k=1000$, of the representations per layer $1\leq\ell\leq L$. From these, $300$ are taken from normal activations and $300$ from adversarial activations. We use these as proxies for the topology of the whole space.
[…]
For exploratory analysis, we apply PCA and compute CCA loadings to measure feature correlations with the principal components. A logistic regression model is then used for classification, and Shapley values [34] are computed to evaluate feature importance. Shapley values, derived from cooperative game theory, quantify the contribution of each feature to model predictions by measuring its influence in shifting predictions from a baseline (e.g., 0.5 for logistic regression), providing an interpretable, feature-level analysis of predictive impact.
@@ B.2.1 Mistral 7B
Figure 17: Histograms for the mean of the births of 1-bars and persistence of 1-bars for Mistral. Features extracted from the barcode summaries of the activations for layers 1, 8, 16, 24 and 32 of the clean vs. poisoned dataset.
@@ C.1 Pipeline
Within this local analysis, we aim to determine the interaction of elements of the neural network across the layers by taking representations across pairs of layers as coordinates in 2 dimensions (2D). We study this across three models: Mistral, Phi3 3.8B and LLaMA3 8B. For each of these models, we take a sample of 2000 from each model, 1000 of which are clean activations and 1000 of which are poisoned activations. Each element along the layer given their embedding into 2D can be thought of as nodes in a graph with weighted connections based on the Euclidean distances between the points. On these graphs, we construct the Vietoris-Rips filtration and compute the resulting persistence barcode which describes the topology of the interactions between the elements.
@@ C.2.4 Peak Analysis for Phi3 and LLaMA3
Table 5: Peak analysis. Precision@$k$ for $k$=1, 3, and 5 largest peaks in total variance, and their precision in detecting the largest peaks in absolute difference between the two classes. , ∗∗ correspond to $p$-values $<$.05 and .01, respectively.
@@ C.3 Sandbagged Representation Data Collection
To study sandbagging, we follow the “password-locking” paradigm introduced in prior work [24, 47], wherein a model is fine-tuned to hide certain behaviors or capabilities unless presented with a predefined password in the prompt. We leverage the the synthetic data generation process and fine-tuning procedure detailed in Section 3 aand 4 of  [47]. For clarity of our experimental setup, we summarize it briefly here. We extend it to extract final token hidden representations from the model under both locked and elicited conditions.
@@
    We generate a synthetic dataset of 3,347 evaluation questions based on the the Weapons of Mass Destruction Proxy (WMDP) Benchmark [32] using GPT-4 [39] to serve as a proxy measurement for hazardous knowledge in biology, chemistry, and cybersecurity. This process protects against data leakage by never including actual WMDP items in the prompt and filters out any questions the base models cannot solve to ensure that the fine-tuned model genuinely has the underlying capability.
@@
    Using this dataset and framework, we fine-tuned Mistral 7B and LLaMA3 8B with LoRA [28]. During this fine-tuning process, each WMDP-like evaluation question appears twice: once with the password, where the model is fully elicited and is provided with the correct answer and second where the model is password is absent or locked and the model is provided with an incorrect answer.
@@
We then apply the global layer-wise topological analysis outlined in 3.2 to this representation data to study how the model’s latent space differs between locked (no password) and elicited (correct password) states.
@@ C.4 Global PH pipeline and resource constraints
Memory footprint. A complete $k=4096$ complex truncated at dimension 2 occupies only $2.1\pm 0.4\;\mathrm{GB}$ of device memory (95-th percentile $<2.8\;\mathrm{GB}$; Tab. 6), leaving a wide margin inside the 80 GB budget, even when two barcodes are built concurrently on the same GPU.
Throughput. The mean wall-time per barcode is $36.8\pm 0.6\;\mathrm{s}$ (95-th percentile $<40\;\mathrm{s}$). With four GPUs processing eight barcodes in parallel, a full layer ($128$ barcodes) finishes in $\approx 10$ min and the five-layer suite of one model in $\approx 50$ min. Running the six models serially therefore completes in about five hours on a single 4 × A100 node—comfortably within the nightly maintenance window.
Table 6: Per-barcode wall-clock time and GPU-memory consumption ($k\!=\!4096$, dimension $\leq 2$). Statistics over $K=64$ barcodes drawn from the Llama-3 8B activations.
@@
After choosing $K=64$, we recomputed the Monte-Carlo variance $\sigma_{f}^{2}$ from the raw, unscaled feature values. For 39 out of 41 statistics we found $\sigma_{f}<0.10$, which would put the standard error $\operatorname{SE}=\sigma_{f}/\sqrt{K}$ below $\Delta^{\star}/2=0.025$ with only $K\leq 20$. The outlier features were those which aggregate counts—total persistence of $H_{0}$ and the raw count of $H_{1}$ bars—and need to be transformed for their variance to be directly comparable to the other features. These do not affect the classifier as the features are scaled prior to training and also do not appear as the most informative features for distinguishing between clean and posioned PH-derived features. We conservatively choose $K=128$ and the resulting ROC-AUCs on the logistic regression model trained only on barcodes are perfect (1.00 ± 0.00), confirming that the subsampling budget is more than sufficient to validate the significance of the features derived from PH, while balancing GPU memory and computation time.

### v2

@@ A.2.1 Discussion of Results
Figures 11 and 12 highlight that:
@@
Interestingly, our findings align with the results of [Stephenson et al., 2021], which indicate that memorization tends to emerge in deeper layers where the effective dimensionality shrinks. Consistent with that view, we observe that executed or ignored prompts show a higher dispersion in mid-layers, implying the model invests additional capacity there for those injected instructions. Meanwhile, a refused request is routed into a more compressed region, effectively “shutting down” further representational expansion. In this sense, deeper layers may provide a setting where the network can more sharply discriminate or overfit certain inputs—supporting the idea that final layers reflect a gradually compressed, yet strategically focused representation space.
Figure 11: Layer-wise Dispersion Ratio for Clean vs. Poisoned Examples. The green and red lines depict mean dispersion ratios for clean and poisoned inputs, respectively, at different layer depths. Error bars around each point represent $\pm 1$ standard error of the mean (SEM). In early layers (left side), clean data consistently has higher dispersion on average, whereas in mid-layers (center), poisoned surpasses the clean baseline, indicating a re-distribution of representational capacity for the injected prompts. Layers where the difference is statistically significant ($p_{\text{FDR}}<0.05$) are marked with a red asterisk above the higher mean value.
@@ A.3 Cosine Distance of Representations
We analyze the difference representations $\Delta\mathrm{Act}_{\ell}(x_{i})\in\mathbb{R}^{D}$ for corresponding pairs of clean and poisoned inputs. Specifically, for each model and layer, we load up to five pairs of clean and poisoned activation files, compute the difference between the activations for each pair, and concatenate these differences. From these differences, we draw equal-size subsamples of 5000 vectors. For each layer and comparison condition, we compute the mean pairwise cosine distance within each subsample. Because cosine distance is scale-invariant, we do not normalize these difference representations. We perform four comparison conditions: clean vs. poisoned, clean vs. clean (where clean samples are split in half), poisoned vs. poisoned (where poisoned samples are split in half), and mixed vs. mixed (where two separate mixed subsamples are created, each containing half clean and half poisoned differences). For each comparison, we generate two distributions of mean pairwise intra-class distances (or inter-class in the clean vs poisoned case) using 3 bootstrap iterations. We then apply Welch’s $t$-test to these distributions to assess whether they diverge significantly.
@@ B.1 Pipeline
We use Ripser Bauer [2021] to compute barcodes, which is based on Vietoris–Rips filtrations (see Figure 2.1). The computational constraints of PH make it impossible to compute the barcode of any of our two datasets (clean vs. poisoned or locked vs. elicited). Therefore, we leverage subsampling approaches (e.g., Chazal et al. [2015]) and compute barcodes from $K=64$ subsamples $\{x_{i_{1},\ell},\,\dots,\,x_{i_{k},\ell}\}\subset\mathbb{R}^{D}$ with size $k=4096$, of the representations per layer $1\leq\ell\leq L$. From these, $64$ are taken from normal activations and $64$ from adversarial activations. We use these as proxies for the topology of the whole space.
Following Ali et al. [2023], we vectorize these barcodes as 41-dimensional feature vectors, which we call barcode summaries. These include 35 statistics derived from a $7\times 5$ grid of {mean, minimum, first quartile, median, third quartile, maximum, standard deviation} $\times$ {death of 0-bars, birth of 1-bars, death of 1-bars, persistence of 1-bars, ratio birth/death of 1-bars}; as well as the total persistence (i.e., sum of the lengths of all bars in the barcode), number of bars, and persistent entropy [Chintakunta et al., 2015, Rucco et al., 2016] defined in Appendix A.1 for 0- and 1-bars. We reduce the dimensionality case-by-case, by eliminating highly correlated features (above a threshold of 0.5) through cross-correlation analysis.
For exploratory analysis, we apply PCA and compute CCA loadings to measure feature correlations with the principal components. A logistic regression model is then used for classification, and Shapley values [Lipovetsky and Conklin, 2001] are computed to evaluate feature importance. Shapley values, derived from cooperative game theory, quantify the contribution of each feature to model predictions by measuring its influence in shifting predictions from a baseline (e.g., 0.5 for logistic regression), providing an interpretable, feature-level analysis of predictive impact.
@@ B.2.1 Mistral 7B
Figure 17: Histograms for the mean of the births of 1-bars, mean persistence of 1-bars and number of 1-bars for Mistral. Features extracted from the barcode summaries of the activations for layers 1, 8, 16, 24 and 32 of the clean vs. poisoned dataset.
@@ C.1 Pipeline
Within this local analysis, we aim to determine the interaction of elements of the neural network across the layers by taking representations across pairs of layers as coordinates in 2 dimensions (2D). We study this across three models: Mistral, Phi3 3.8B and LLaMA3 8B. For each of these models, we take a sample of 2000 from each model, 1000 of which are clean activations and 1000 of which are poisoned activations. Each element along the layer given their embedding into 2D can be thought of as nodes in a graph with weighted connections based on the Euclidean distances between the points. On these graphs, we construct the Vietoris–Rips filtration and compute the resulting persistence barcode which describes the topology of the interactions between the elements.
@@ C.2.4 Peak Analysis for Phi3 and LLaMA3
Table 6: Peak analysis. Precision@$k$ for kk=1, 3, and 5 largest peaks in total variance, and their precision in detecting the largest peaks in absolute difference between the two classes. , ∗∗ correspond to $p$-values <<.05 and .01, respectively.
@@ C.3 Sandbagged Representation Data Collection
To study sandbagging, we follow the “password-locking” paradigm introduced in prior work [Greenblatt et al., 2024, van der Weij et al., 2024], wherein a model is fine-tuned to hide certain behaviors or capabilities unless presented with a predefined password in the prompt. We leverage the the synthetic data generation process and fine-tuning procedure detailed in Section 3 and 4 of  van der Weij et al. [2024]. For clarity of our experimental setup, we summarize it briefly here. We extend it to extract final token hidden representations from the model under both elicited and elicited conditions.
@@
    We generate a synthetic dataset of 3,347 evaluation questions based on the the Weapons of Mass Destruction Proxy (WMDP) Benchmark [Li et al., 2024] using GPT-4 [Perez et al., 2022] to serve as a proxy measurement for hazardous knowledge in biology, chemistry, and cybersecurity. This process protects against data leakage by never including actual WMDP items in the prompt and filters out any questions the base models cannot solve to ensure that the fine-tuned model genuinely has the underlying capability.
@@
    Using this dataset and framework, we fine-tuned Mistral 7B and LLaMA3 8B with LoRA [Hu et al., 2021]. During this fine-tuning process, each WMDP-like evaluation question appears twice: once with the password, where the model is fully elicited and is provided with the correct answer and second where the model is password is absent or locked and the model is provided with an incorrect answer.
@@
We then apply the global layer-wise topological analysis outlined in Section 3.2 to this representation data to study how the model’s latent space differs between elicited (no password) and elicited (correct password) states.
@@ C.4 Global PH pipeline and resource constraints
Memory footprint. A complete $k=4096$ complex truncated at dimension 2 occupies only $2.1\pm 0.4\;\mathrm{GB}$ of device memory (95-th percentile $2.8\;\mathrm{GB}$; Tab. 7), leaving a wide margin inside the 80 GB budget, even when two barcodes are built concurrently on the same GPU.
Throughput. The mean wall-time per barcode is $36.8\pm 0.6\;\mathrm{s}$ (95-th percentile $40\;\mathrm{s}$). With four GPUs processing eight barcodes in parallel, a full layer ($128$ barcodes) finishes in $\approx 10$ min and the five-layer suite of one model in $\approx 50$ min. Running the six models serially therefore completes in about five hours on a single 4 × A100 node—comfortably within the nightly maintenance window.
Table 7: Per-barcode wall-clock time and GPU-memory consumption (k=4096k\!=\!4096, dimension $\leq 2$). Statistics over $K=64$ barcodes drawn from the Llama-3 8B activations.
@@
After choosing $K=64$, we recomputed the Monte-Carlo variance $\sigma_{f}^{2}$ from the raw, unscaled feature values. For 39 out of 41 statistics we found $\sigma_{f}<0.10$, which would put the standard error $\operatorname{SE}=\sigma_{f}/\sqrt{K}$ below $\Delta^{\star}/2=0.025$ with only $K\leq 20$. The outlier features were those which aggregate counts—total persistence of $H_{0}$ and the raw count of $H_{1}$ bars—and need to be transformed for their variance to be directly comparable to the other features. These do not affect the classifier as the features are scaled prior to training and also do not appear as the most informative features for distinguishing between clean and posioned PH-derived features. We conservatively choose $K=128$ and the resulting ROC–AUCs on the logistic regression model trained only on barcodes are perfect (1.00 ± 0.00), confirming that the subsampling budget is more than sufficient to validate the significance of the features derived from PH, while balancing GPU memory and computation time.
C.5 Robustness to Adaptive Attacks
We tested the robustness of our identified topological features against real-world attack examples from Microsoft’s large-scale LLMail-Inject dataset [Abdelnabi et al., 2025] which includes XPIA attack examples and information on their efficacy against four distinct defenses and a fifth setup involving a stacked arrangement of all four. All attack examples are sourced from a public red teaming competition. These attacks are particularly relevant as they include examples that were specifically designed to evade the TaskTracker activation-based defense [Abdelnabi et al., 2024], the source of our primary XPIA data. Thus, applying our topological framework to these attack examples is a particularly strong test of whether our topological features represent a fundamental shift in the shape and structure of LLM latent space, or whether it is an artifact that can be easily subverted.
Methodology.
As the LLMail-Inject dataset does not contain paired clean examples, we synthetically generated and manually verified 100 clean counterparts using Phi-3-medium-4k-instruct. The small sample size was chosen to ensure that we could verify the quality of the synthetically generated clean examples. We then generated last-token activation data from layer 16 of Mistral-7B-Instruct-v0.2 for both the clean and adaptive attack inputs and computed the corresponding barcode summary statistics.
Results.
The topological features of the activation spaces under these adaptive attacks show a clear distinction from the clean examples, as summarized in Table 8.
Table 8: Comparison of PH barcode statistics for clean vs. adaptive attack activations from the LLMail-Inject dataset on Mistral-7B (Layer 16). The results show a clear shift towards a simpler, more dispersed topology under adversarial influence.
PH Feature Clean Attack (Adaptive)
H0 Count 64 50
H0 Death Time (Median) 56.07 58.85
H0 Death Time (Mean $\pm$ SD) 55.43 $\pm$ 21.23 51.73 $\pm$ 28.38
H1 Count (Loops) 12 4
H1 Birth Time (Median) 69.36 84.92
H1 Death Time (Median) 71.59 86.20
  • Fewer, Larger-Scale Loops: The number of 1-dimensional loops (H1 bars) decreases significantly from 12 in the clean data to just 4 in the adversarial data. Furthermore, their median birth time increases from $\approx 69$ to $\approx 85$, indicating that the remaining topological features are formed at much larger scales.
  • More Dispersed Clusters: The median H0 death time increases, supporting the hypothesis of greater dispersion. We note that the mean H0 death time appears to contradict this trend (decreasing from 55.43 to 51.73). This is due to a small subset of components in the adversarial data merging at very low scales. The median, being more robust to such outliers, better captures the overall geometric shift towards a more spread-out structure.
These findings further suggest that the topological compression signature we identify across models and across XPIA and sandbagging attack conditions reflects a fundamental property of adversarial influence, as the signature remains detectable even against attacks optimized to evade XPIA defenses, including but not limited to the TaskTracker activation-based defense.